OpenStudy (anonymous):

Simplify each expression. Show work. 1/4(8x-4)-1/5(20x-6y)

OpenStudy (anonymous):

Hint:\[\frac{ a }{ b } \pm \frac{ c }{ d } \implies \frac{ ad \pm bc }{ bd }\]

OpenStudy (anonymous):

Besides that im not using that formula

OpenStudy (anonymous):

I'll give medal for those who help me solve this problem on paper

JoelTheBoss (joel_the_boss):

\[\huge{\color{red}{\star} ^\color{blue}{{\bigstar} ^\color{green}{{\star} ^\color{magenta}{{\bigstar }}}}}\color{cyan}{Welcome ~~To~~Openstudy!!}~^{\color{red}{\star} ^\color{blue}{{\bigstar} ^\color{green}{{\star} ^\color{magenta}{{\bigstar }}}}} \] \[\frac{1}{4}(8x-4)-\frac{1}{5}(20x-6y)\]Is that what it should ook like? @clogger5

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I need help solving it on paper. Medals to those who help solve problem.